OpenStudy (anonymous):

Express each repeating decimal as a fraction. 0.7, 0.42, and 0.16

OpenStudy (anonymous):

these do not repeat. do you mean .777... ?

OpenStudy (anonymous):

these do not repeat. do you mean .777... ?

OpenStudy (anonymous):

otherwise \[.7=\frac{7}{10}\] whereas \[.777...=\frac{7}{9}\]

OpenStudy (anonymous):

0.42 repeating = 42/99?

OpenStudy (anonymous):

and \[.42=\frac{42}{100}=\frac{21}{50}\] whereas \[.424242...=\frac{42}{99}=\frac{14}{33}\]

OpenStudy (anonymous):

I dont know how to write it but there is a line over the last number in each

OpenStudy (anonymous):

ok then answers are correct

OpenStudy (anonymous):

\[\overline{.7}\] means .7777... that is the sevens repeat, and the fraction is \[\frac{7}{9}\]

OpenStudy (anonymous):

likewise \[\overline{.42}\] means .424242... and the fraction is \[\frac{42}{99}=\frac{14}{33}\]

OpenStudy (anonymous):

and 0.16?

OpenStudy (anonymous):

line over 6

OpenStudy (anonymous):

0.161616.. = 16/99 0.16 = 16/100 = 4/25

OpenStudy (anonymous):

if it means .166666... then different answer

OpenStudy (anonymous):

\[.1\bar{6}\]?

OpenStudy (anonymous):

yes but 0 before the point

OpenStudy (anonymous):

0 is unimportant.

OpenStudy (anonymous):

do this: put \[x=0.1666...\] \[10x=1.6666...\] \[9x=10x-x=1.5\] \[x=\frac{1.5}{9}=\frac{15}{90}=\frac{1}{6}\]

OpenStudy (anonymous):

so the final answer is 1/6

OpenStudy (anonymous):

?